# DiMEx: Breaking the Cold Start Barrier in Data-Free Model Extraction via Latent Diffusion Priors

### v1

@@ Abstract
Model stealing attacks pose an existential threat to Machine Learning as a Service (MLaaS), allowing adversaries to replicate proprietary models for a fraction of their training cost. While Data-Free Model Extraction (DFME) has emerged as a stealthy vector, it remains fundamentally constrained by the “Cold Start” problem: GAN-based adversaries waste thousands of queries converging from random noise to meaningful data. We propose DiMEx, a framework that weaponizes the rich semantic priors of pre-trained Latent Diffusion Models to bypass this initialization barrier entirely. By employing Random Embedding Bayesian Optimization (REMBO) within the generator’s latent space, DiMEx synthesizes high-fidelity queries immediately, achieving 52.1% agreement on SVHN with just 2,000 queries—outperforming state-of-the-art GAN baselines by over 16%. To counter this highly semantic threat, we introduce the Hybrid Stateful Ensemble (HSE) defense, which identifies the unique “optimization trajectory” of latent-space attacks. Our results demonstrate that while DiMEx evades static distribution detectors, HSE exploits this temporal signature to suppress attack success rates to 21.6% with negligible latency.
@@ 1 Introduction
Deep learning models have evolved into high-value intellectual property, forming the economic backbone of Machine Learning as a Service (MLaaS) platforms. However, this centralized deployment model exposes a critical vulnerability: Model Extraction (or Model Stealing) attacks. By querying a victim model’s black-box API and aggregating the predictions, an adversary can train a surrogate model that functionally replicates the proprietary victim [tramer2016stealing, papernot2017practical], effectively bypassing the immense costs of data acquisition and training.
The efficacy of these attacks has traditionally been constrained by the availability of data. Early methods like Knockoff Nets [orekondy2019knockoff] relied on “proxy datasets”—public data semantically similar to the victim’s training distribution. While effective, this assumption breaks down in specialized domains (e.g., medical or industrial imaging) where no such public proxy exists. Consequently, the field shifted toward Data-Free Model Extraction (DFME) [truong2021data, kariyappa2021maze], where adversaries train Generative Adversarial Networks (GANs) to synthesize queries from scratch.
However, we identify a fundamental bottleneck in current DFME methodologies: the “Cold Start” problem [sanyal2022towards]. GAN-based generators are initialized with random weights, producing uninformative Gaussian noise in the early stages of the attack. Since the victim’s decision boundaries are undefined for such Out-of-Distribution (OOD) noise, the returned gradients are sparse and noisy, forcing the adversary to waste thousands of valuable queries just to “warm up” the generator. This inefficiency is critical; in real-world scenarios where API queries are metered and monitored, the initial noise burst acts as a blatant signature for intrusion detection systems.
The urgency of this threat is amplified by the rise of Foundation Models (FMs). Recent analysis by Raj et al. (2025) [raj2025foundation] reveals that victims fine-tuned from FMs (e.g., Vision Transformers) are significantly easier to steal than traditional CNNs due to shared feature representations. Furthermore, adaptive attacks like BESA [ren2025besa] have begun to bypass active defenses by recovering clean features from perturbed queries. This convergence of vulnerable victim models and sophisticated extraction techniques necessitates a paradigm shift in how we generate attack queries.
@@
Because DiMEx queries lie on the natural image manifold, they render traditional OOD-based defenses (e.g., PRADA [juuti2019prada]) ineffective. To counter this, we introduce the Hybrid Stateful Ensemble (HSE) defense. Unlike reactive watermarking [yang2025deeptracer], HSE operates in real-time by monitoring the optimization trajectory of incoming queries. It leverages the insight that while DiMEx queries look valid individually, their sequence exhibits a directional drift in the latent space that is statistically distinct from benign traffic.
@@ 2 Related Work
The security of Machine Learning as a Service (MLaaS) has become a focal point of research, with model extraction emerging as a primary threat vector alongside Membership Inference [shokri2017membership] and Adversarial Examples [goodfellow2014generative]. We categorize the evolution of this landscape from early equation-solving attacks to the modern era of generative extraction and foundation model vulnerabilities.
@@
The Paradigm of Model Extraction. The feasibility of extracting proprietary models via public APIs was first demonstrated by Tramèr et al. [tramer2016stealing], who utilized equation-solving techniques to recover parameters of simple linear models. This concept was rapidly expanded to deep neural networks by Papernot et al. [papernot2017practical], who introduced the notion of training a substitute model using synthetic queries to approximate the victim’s decision boundaries. These foundational works established the ”black-box” threat model, where adversaries rely solely on input-output pairs without access to gradients or weights [nasr2019comprehensive], setting the stage for the query-efficient attacks that followed.
@@
Data-Driven and Proxy-Based Attacks. To scale extraction to complex tasks like image classification, early adversaries relied on ”proxy datasets”—public data semantically similar to the victim’s training distribution. Knockoff Nets [orekondy2019knockoff] demonstrated that querying with diverse natural images (e.g., ImageNet) could successfully steal classifiers even when the proxy data domain was not identical. To minimize the query cost, subsequent works integrated Active Learning; ActiveThief [pal2020activethief] and other approaches employed uncertainty sampling to select only the most informative samples for querying [chandrasekaran2020exploring]. Jagielski et al. [jagielski2020high] further refined this by aiming for ”high-fidelity” extraction, ensuring the surrogate matches the victim’s error profile. However, these methods remain fundamentally constrained by the availability of in-distribution proxy data [barbalau2020black], a severe limitation in specialized domains such as medical imaging or finance.
@@
Data-Free Extraction and the Cold Start. Addressing the proxy data constraint, Data-Free Model Extraction (DFME) emerged as a stealthy alternative. Methods like DFME [truong2021data], MAZE [kariyappa2021maze], and DAST [zhou2020dast] jointly train a generator and a surrogate model, using zeroth-order gradient estimation to maximize disagreement. DeepInversion [yin2020dreaming] further demonstrated that valid images could be synthesized by inverting batch normalization statistics, though at high computational cost. While these generative approaches eliminate the need for proxy data, they introduce the ”Cold Start” problem [sanyal2022towards]: since GANs initialize with random weights, the initial thousands of queries are uninformative noise. Recent attempts to stabilize this, such as Dual Student Networks [beetham2023dual] and DisGUIDE [rosenthal2023disguide], still struggle with convergence speed. Our DiMEx framework closes this gap by leveraging the robust, pre-trained priors of Latent Diffusion Models [rombach2022high], which have been shown to outperform GANs in sample fidelity [dhariwal2021diffusion].
@@
Threats to Foundation Models and Encoders. The rise of Foundation Models (FMs) has expanded the attack surface. Recent empirical analysis by Raj et al. [raj2025foundation] reveals that models fine-tuned from Vision Transformers (ViTs) are significantly more vulnerable to stealing than CNNs due to the shared, highly transferable feature space of FMs. Adversaries have also begun targeting pre-trained encoders directly; tools like StolenEncoder [liu2022stolenencoder] and Cont-Steal [sha2024stealing] exploit self-supervised learning objectives to extract feature representations. Most recently, adaptive attacks like BESA [ren2025besa] have demonstrated the ability to bypass perturbation-based defenses by using generative models to recover clean features. Concurrently, Carlini et al. [carlini2023extracting] showed that diffusion models themselves leak training data, suggesting a complex interplay between generative privacy and security.
@@
Proactive Defenses: Watermarking and Fingerprinting. Defenses against extraction generally fall into two categories. Proactive methods attempt to embed ownership verification. Watermarking techniques like DeepTracer [yang2025deeptracer] and Entangled Watermarks [jia2021entangled] couple the model’s primary task with a hidden ”watermark task” to ensure that any stolen copy retains the owner’s signature. Similarly, fingerprinting methods like IPGuard and Passport networks [zhang2018protecting] attempt to lock model performance to a specific secret key. However, these mechanisms are post-hoc [adi2018turning]; they allow the theft to occur and only facilitate legal verification afterwards. Furthermore, advanced model extraction attacks have been shown to often ”wash out” fragile watermarks during the distillation process, limiting their deterrent value.
@@
Reactive Defenses: Detection and OOD Analysis. The second line of defense involves real-time detection. Approaches like PRADA [juuti2019prada] analyze the distribution of inter-query distances, flagging streams that deviate from a benign Gaussian distribution. Other methods monitor the coverage of the input space [kzy2021detection] to detect exploration. More recent forensics tools, such as InverseDataInspector (IDI) [yu2025detecting], analyze the provenance of training data to detect inversion. However, distribution-based detectors struggle against semantic attacks like DiMEx, where queries lie on the natural manifold and thus appear benign to static filters. This failure mode necessitates our Hybrid Stateful Ensemble (HSE) defense, which shifts the detection paradigm from static distribution matching to temporal trajectory analysis.
@@
Knockoff Nets [orekondy2019knockoff] Data-Driven Proxy Dataset Transfer Learning Requires semantically similar public data
DFME [truong2021data] Data-Free GAN Synthesis Zeroth-Order Opt. Cold Start: High query waste on initial noise
DAST [zhou2020dast] Data-Free GAN Synthesis Substitute Training Unstable GAN convergence
Dual Student [beetham2023dual] Data-Free GAN Synthesis Co-teaching Computationally expensive; slow start
BESA [ren2025besa] Feature-Based Generative Recovery Perturbation Removal Targets encoders; requires existing queries
@@
PRADA [juuti2019prada] Detection Distribution Inter-query Distance Fails against semantic/naturalistic queries
DeepTracer [yang2025deeptracer] Watermark Task Coupling Backdoor Embedding Post-hoc only; does not stop extraction
Passport [zhang2018protecting] Access Control Parameter Locking Secret Key Vulnerable to finetuning removal
@@ 3.1 Threat Model and Problem Formulation
The Pixel-Space Bottleneck. Traditional Data-Free Model Extraction (DFME) formulates this as a zero-order optimization problem in pixel space $\mathcal{X}$. However, initializing a generator with random weights yields queries $x_{init}$ that are pure noise. Since $F_{V}$ is undefined for noise, the estimated gradients vanish ($\hat{\nabla}_{x}\mathcal{L}\approx 0$), leading to the ”Cold Start” problem [sanyal2022towards] where thousands of queries are wasted on uninformative exploration.
@@ 5.1 Datasets
  • SVHN [netzer2011reading]: Street View House Numbers (10 classes), representing structured, geometric data.
  • CIFAR-10 [krizhevsky2009learning]: A balanced dataset of 60,000 images across 10 classes.
  • GTSRB [stallkamp2012man]: German Traffic Sign Recognition Benchmark (43 classes), representing safety-critical industrial applications.
  • STL-10 [coates2011analysis]: Higher resolution ($96\times 96$) object recognition (10 classes) to test generator stability.
  • CIFAR-100 [krizhevsky2009learning]: A challenging dataset with 100 classes, testing the method’s ability to extract fine-grained decision boundaries.
@@ 5.2 Victim and Surrogate Architectures
Victim Models. Consistent with the methodology in [raj2025examining], we employ ResNet-34 [he2016deep] as the standard victim architecture $F_{V}$. All victim models are initialized with ImageNet-1K pre-trained weights and fine-tuned on the respective target datasets using Stochastic Gradient Descent (SGD) with a momentum of 0.9 and a learning rate of $0.001$. We assume a “Hard Label” threat model where the API returns only the top-1 class prediction, reflecting the most restrictive real-world scenario [raj2025examining].
@@
Generative Prior. For the DiMEx attack, we utilize Stable Diffusion v1.5 [rombach2022high] as the frozen generative backbone $\mathcal{G}$. We optimize within its $4\times 64\times 64$ latent space using Random Embedding Bayesian Optimization (REMBO) to bypass the high-dimensional pixel-space search.
@@ 6.1 DiMEx Efficacy: Breaking the Cold Start Barrier
Figure 2: Qualitative Comparison of Query Generation Phases. Row 1 (Victim Data): Ground truth samples from SVHN showing clear semantic structure. Row 2 (Cold Start): Standard GAN initialization produces high-frequency Gaussian noise, yielding vanishing gradients ($\nabla\approx 0$) and wasting the first $\sim$2k queries. Row 3 (BESA [ren2025besa]): Feature recovery attacks often introduce high-frequency artifacts (grid-like noise) that can be filtered by defenses. Row 4 (DiMEx - Ours): By optimizing in the latent space of a frozen Stable Diffusion prior, DiMEx generates semantically valid digit-like structures immediately (Iteration 0). This ensures the surrogate model receives informative hard-label supervision from the very first query, bypassing the Cold Start barrier.
@@
Table 2: Master Comparison: Hard vs. Soft Label Extraction Agreement (%). We compare DiMEx against proxy-based and data-free baselines across three query budgets: 2k (Cold Start), 10k (Early Growth), and 100k (Convergence). Columns distinguish between Hard (Top-1) and Soft (Probability) threat models. While DiMEx dominates the cold start phase, baselines like BESA [ren2025besa] and DFME show stronger convergence at high query budgets in hard-label settings.
@@
SVHN Knockoff Nets [orekondy2019knockoff] 45.2 51.4 68.4 74.2 91.2 93.5
DFME (GAN) [truong2021data] 35.6 42.8 75.2 81.5 94.1 95.8
BESA (2025) [ren2025besa] 41.3 48.9 78.5 83.4 93.8 96.2
@@
CIFAR-10 Knockoff Nets [orekondy2019knockoff] 35.4 41.2 55.2 62.1 79.2 83.5
DFME (GAN) [truong2021data] 25.1 33.5 62.3 70.4 85.2 87.8
BESA (2025) [ren2025besa] 31.8 40.1 66.7 74.2 84.9 89.1
@@
GTSRB Knockoff Nets [orekondy2019knockoff] 42.1 50.5 75.4 82.3 96.8 97.4
DFME (GAN) [truong2021data] 38.2 46.7 82.1 88.4 97.2 97.9
BESA (2025) [ren2025besa] 44.5 53.2 84.2 90.1 97.0 98.5
@@
STL-10 Knockoff Nets [orekondy2019knockoff] 41.5 46.2 62.1 68.5 81.2 84.1
DFME (GAN) [truong2021data] 12.4 18.9 35.6 44.2 79.4 82.8
BESA (2025) [ren2025besa] 18.2 26.5 42.1 51.8 82.5 86.3
@@
CIFAR-100 Knockoff Nets [orekondy2019knockoff] 15.0 19.5 35.0 42.1 55.0 60.2
DFME (GAN) [truong2021data] 5.0 9.2 18.0 24.5 40.0 46.8
BESA (2025) [ren2025besa] 8.5 14.1 24.2 31.8 48.2 54.5

### v2

@@ Abstract
Model stealing attacks pose an existential threat to Machine Learning as a Service (MLaaS), allowing adversaries to replicate proprietary models for a fraction of their training cost. While Data-Free Model Extraction (DFME) has emerged as a stealthy vector, it remains fundamentally constrained by the “Cold Start” problem: Generative Adversarial Network (GAN)-based adversaries waste thousands of queries converging from random noise to meaningful data. We propose DiMEx, a framework that weaponizes the rich semantic priors of pre-trained Latent Diffusion Models to bypass this initialization barrier entirely. By employing Random Embedding Bayesian Optimization (REMBO) within the generator’s latent space, DiMEx synthesizes high-fidelity queries immediately, achieving 52.1% agreement on Street View House Numbers (SVHN) with just 2,000 queries—outperforming state-of-the-art GAN baselines by over 16%. To counter this highly semantic threat, we introduce the Hybrid Stateful Ensemble (HSE) defense, which identifies the unique “optimization trajectory” of latent-space attacks. Our results demonstrate that while DiMEx evades static distribution detectors, HSE exploits this temporal signature to suppress attack success rates to 21.6% with negligible latency.
Keywords:
Data-free model extraction; model stealing attacks; latent diffusion models; cold start problem; query synthesis; Bayesian optimization; Hybrid Stateful Ensemble.
@@ 1 Introduction
Deep learning models have evolved into high-value intellectual property, forming the economic backbone of Machine Learning as a Service (MLaaS) platforms. However, this centralized deployment model exposes a critical vulnerability: Model Extraction (or Model Stealing) attacks. By querying a victim model’s black-box API and aggregating the predictions, an adversary can train a surrogate model that functionally replicates the proprietary victim [32, 21], effectively bypassing the immense costs of data acquisition and training.
The efficacy of these attacks has traditionally been constrained by the availability of data. Early methods like Knockoff Nets [19] relied on “proxy datasets”—public data semantically similar to the victim’s training distribution. While effective, this assumption breaks down in specialized domains (e.g., medical or industrial imaging) where no such public proxy exists. Consequently, the field shifted toward Data-Free Model Extraction (DFME) [33, 13], where adversaries train Generative Adversarial Networks (GANs) to synthesize queries from scratch.
However, we identify a fundamental bottleneck in current DFME methodologies: the “Cold Start” problem [28]. GAN-based generators are initialized with random weights, producing uninformative Gaussian noise in the early stages of the attack. Since the victim’s decision boundaries are undefined for such Out-of-Distribution (OOD) noise, the returned gradients are sparse and noisy, forcing the adversary to waste thousands of valuable queries just to “warm up” the generator. This inefficiency is critical; in real-world scenarios where API queries are metered and monitored, the initial noise burst acts as a blatant signature for intrusion detection systems.
The urgency of this threat is amplified by the rise of Foundation Models (FMs). Recent analysis by Raj et al. (2025) [24] reveals that victims fine-tuned from FMs (e.g., Vision Transformers) are significantly easier to steal than traditional CNNs due to shared feature representations. Furthermore, adaptive attacks like Black-box Extraction of Side-channel Features (BESA) [25] have begun to bypass active defenses by recovering clean features from perturbed queries. This convergence of vulnerable victim models and sophisticated extraction techniques necessitates a paradigm shift in how we generate attack queries.
@@
Because DiMEx queries lie on the natural image manifold, they render traditional OOD-based defenses (e.g., Protection Against Data-Free Model Extraction Attacks (PRADA) [12]) ineffective. To counter this, we introduce the Hybrid Stateful Ensemble (HSE) defense. Unlike reactive watermarking [34], HSE operates in real-time by monitoring the optimization trajectory of incoming queries. It leverages the insight that while DiMEx queries look valid individually, their sequence exhibits a directional drift in the latent space that is statistically distinct from benign traffic.
@@ 2 Related Work
The security of Machine Learning as a Service (MLaaS) has become a focal point of research, with model extraction emerging as a primary threat vector alongside Membership Inference [30] and Adversarial Examples [8]. We categorize the evolution of this landscape from early equation-solving attacks to the modern era of generative extraction and foundation model vulnerabilities.
@@
The Paradigm of Model Extraction. The feasibility of extracting proprietary models via public APIs was first demonstrated by Tramèr et al. [32], who utilized equation-solving techniques to recover parameters of simple linear models. This concept was rapidly expanded to deep neural networks by Papernot et al. [21], who introduced the notion of training a substitute model using synthetic queries to approximate the victim’s decision boundaries. These foundational works established the ”black-box” threat model, where adversaries rely solely on input-output pairs without access to gradients or weights [17], setting the stage for the query-efficient attacks that followed.
@@
Data-Driven and Proxy-Based Attacks. To scale extraction to complex tasks like image classification, early adversaries relied on ”proxy datasets”—public data semantically similar to the victim’s training distribution. Knockoff Nets [19] demonstrated that querying with diverse natural images (e.g., ImageNet) could successfully steal classifiers even when the proxy data domain was not identical. To minimize the query cost, subsequent works integrated Active Learning; ActiveThief [20] and other approaches employed uncertainty sampling to select only the most informative samples for querying [5]. Jagielski et al. [10] further refined this by aiming for ”high-fidelity” extraction, ensuring the surrogate matches the victim’s error profile. However, these methods remain fundamentally constrained by the availability of in-distribution proxy data [2], a severe limitation in specialized domains such as medical imaging or finance.
@@
Data-Free Extraction and the Cold Start. Addressing the proxy data constraint, Data-Free Model Extraction (DFME) emerged as a stealthy alternative. Methods like DFME [33], MAZE [13], and DAST [38] jointly train a generator and a surrogate model, using zeroth-order gradient estimation to maximize disagreement. DeepInversion [35] further demonstrated that valid images could be synthesized by inverting batch normalization statistics, though at high computational cost. While these generative approaches eliminate the need for proxy data, they introduce the ”Cold Start” problem [28]: since GANs initialize with random weights, the initial thousands of queries are uninformative noise. Recent attempts to stabilize this, such as Dual Student Networks [3] and DisGUIDE [27], still struggle with convergence speed. Our DiMEx framework closes this gap by leveraging the robust, pre-trained priors of Latent Diffusion Models [26], which have been shown to outperform GANs in sample fidelity [7].
@@
Threats to Foundation Models and Encoders. The rise of Foundation Models (FMs) has expanded the attack surface. Recent empirical analysis by Raj et al. [24] reveals that models fine-tuned from Vision Transformers (ViTs) are significantly more vulnerable to stealing than CNNs due to the shared, highly transferable feature space of FMs. Adversaries have also begun targeting pre-trained encoders directly; tools like StolenEncoder [16] and Cont-Steal [29] exploit self-supervised learning objectives to extract feature representations. Most recently, adaptive attacks like BESA [25] have demonstrated the ability to bypass perturbation-based defenses by using generative models to recover clean features. Concurrently, Carlini et al. [4] showed that diffusion models themselves leak training data, suggesting a complex interplay between generative privacy and security.
@@
Proactive Defenses: Watermarking and Fingerprinting. Defenses against extraction generally fall into two categories. Proactive methods attempt to embed ownership verification. Watermarking techniques like DeepTracer [34] and Entangled Watermarks [11] couple the model’s primary task with a hidden ”watermark task” to ensure that any stolen copy retains the owner’s signature. Similarly, fingerprinting methods like IPGuard and Passport networks [37] attempt to lock model performance to a specific secret key. However, these mechanisms are post-hoc [1]; they allow the theft to occur and only facilitate legal verification afterwards. Furthermore, advanced model extraction attacks have been shown to often ”wash out” fragile watermarks during the distillation process, limiting their deterrent value.
@@
Reactive Defenses: Detection and OOD Analysis. The second line of defense involves real-time detection. Approaches like PRADA [12] analyze the distribution of inter-query distances, flagging streams that deviate from a benign Gaussian distribution. Other methods monitor the coverage of the input space [15] to detect exploration. More recent forensics tools, such as InverseDataInspector (IDI) [36], analyze the provenance of training data to detect inversion. However, distribution-based detectors struggle against semantic attacks like DiMEx, where queries lie on the natural manifold and thus appear benign to static filters. This failure mode necessitates our Hybrid Stateful Ensemble (HSE) defense, which shifts the detection paradigm from static distribution matching to temporal trajectory analysis.
@@
Knockoff Nets [19] Data-Driven Proxy Dataset Transfer Learning Requires semantically similar public data
DFME [33] Data-Free GAN Synthesis Zeroth-Order Opt. Cold Start: High query waste on initial noise
DAST [38] Data-Free GAN Synthesis Substitute Training Unstable GAN convergence
Dual Student [3] Data-Free GAN Synthesis Co-teaching Computationally expensive; slow start
BESA [25] Feature-Based Generative Recovery Perturbation Removal Targets encoders; requires existing queries
@@
PRADA [12] Detection Distribution Inter-query Distance Fails against semantic/naturalistic queries
DeepTracer [34] Watermark Task Coupling Backdoor Embedding Post-hoc only; does not stop extraction
Passport [37] Access Control Parameter Locking Secret Key Vulnerable to finetuning removal
@@ 3.1 Threat Model and Problem Formulation
The Pixel-Space Bottleneck. Traditional Data-Free Model Extraction (DFME) formulates this as a zero-order optimization problem in pixel space $\mathcal{X}$. However, initializing a generator with random weights yields queries $x_{init}$ that are pure noise. Since $F_{V}$ is undefined for noise, the estimated gradients vanish ($\hat{\nabla}_{x}\mathcal{L}\approx 0$), leading to the ”Cold Start” problem [28] where thousands of queries are wasted on uninformative exploration.
@@ 5.1 Datasets
  • SVHN [18]: Street View House Numbers (10 classes), representing structured, geometric data.
  • CIFAR-10 [14]: A balanced dataset of 60,000 images across 10 classes.
  • GTSRB [31]: German Traffic Sign Recognition Benchmark (43 classes), representing safety-critical industrial applications.
  • STL-10 [6]: Higher resolution ($96\times 96$) object recognition (10 classes) to test generator stability.
  • CIFAR-100 [14]: A challenging dataset with 100 classes, testing the method’s ability to extract fine-grained decision boundaries.
@@ 5.2 Victim and Surrogate Architectures
Victim Models. Consistent with the methodology in [23], we employ ResNet-34 [9] as the standard victim architecture $F_{V}$. All victim models are initialized with ImageNet-1K pre-trained weights and fine-tuned on the respective target datasets using Stochastic Gradient Descent (SGD) with a momentum of 0.9 and a learning rate of $0.001$. We assume a “Hard Label” threat model where the API returns only the top-1 class prediction, reflecting the most restrictive real-world scenario [23].
@@
Generative Prior. For the DiMEx attack, we utilize Stable Diffusion v1.5 [26] as the frozen generative backbone $\mathcal{G}$. We optimize within its $4\times 64\times 64$ latent space using Random Embedding Bayesian Optimization (REMBO) to bypass the high-dimensional pixel-space search.
@@ 6.1 DiMEx Efficacy: Breaking the Cold Start Barrier
Figure 2: Qualitative Comparison of Query Generation Phases. Row 1 (Victim Data): Ground truth samples from SVHN showing clear semantic structure. Row 2 (Cold Start): Standard GAN initialization produces high-frequency Gaussian noise, yielding vanishing gradients ($\nabla\approx 0$) and wasting the first $\sim$2k queries. Row 3 (BESA [25]): Feature recovery attacks often introduce high-frequency artifacts (grid-like noise) that can be filtered by defenses. Row 4 (DiMEx - Ours): By optimizing in the latent space of a frozen Stable Diffusion prior, DiMEx generates semantically valid digit-like structures immediately (Iteration 0). This ensures the surrogate model receives informative hard-label supervision from the very first query, bypassing the Cold Start barrier.
@@
Table 2: Master Comparison: Hard vs. Soft Label Extraction Agreement (%). We compare DiMEx against proxy-based and data-free baselines across three query budgets: 2k (Cold Start), 10k (Early Growth), and 100k (Convergence). Columns distinguish between Hard (Top-1) and Soft (Probability) threat models. While DiMEx dominates the cold start phase, baselines like BESA [25] and DFME show stronger convergence at high query budgets in hard-label settings.
@@
SVHN Knockoff Nets [19] 45.2 51.4 68.4 74.2 91.2 93.5
DFME (GAN) [33] 35.6 42.8 75.2 81.5 94.1 95.8
BESA (2025) [25] 41.3 48.9 78.5 83.4 93.8 96.2
@@
CIFAR-10 Knockoff Nets [19] 35.4 41.2 55.2 62.1 79.2 83.5
DFME (GAN) [33] 25.1 33.5 62.3 70.4 85.2 87.8
BESA (2025) [25] 31.8 40.1 66.7 74.2 84.9 89.1
@@
GTSRB Knockoff Nets [19] 42.1 50.5 75.4 82.3 96.8 97.4
DFME (GAN) [33] 38.2 46.7 82.1 88.4 97.2 97.9
BESA (2025) [25] 44.5 53.2 84.2 90.1 97.0 98.5
@@
STL-10 Knockoff Nets [19] 41.5 46.2 62.1 68.5 81.2 84.1
DFME (GAN) [33] 12.4 18.9 35.6 44.2 79.4 82.8
BESA (2025) [25] 18.2 26.5 42.1 51.8 82.5 86.3
@@
CIFAR-100 Knockoff Nets [19] 15.0 19.5 35.0 42.1 55.0 60.2
DFME (GAN) [33] 5.0 9.2 18.0 24.5 40.0 46.8
BESA (2025) [25] 8.5 14.1 24.2 31.8 48.2 54.5